CALCULATE THE SLOPE, THE X & Y INTERCEPT , THEIR CORRESPONDING POINTS OF ORIGIN AND THEN GRAPH THE FOLLOWING ORDERED PAIRS. 8X-4Y=24
Answer by JBarnum(2146)

first take the 4 out
2x-y=6
2x=6+y
2x-6=y
m=2
y-int=-6
x-int=3 (x=(1/2)y+3)
(0,-6)
(3,0)
you should be able to graph it.